# Score Forgetting Distillation: A Swift, Data-Free Method for Machine Unlearning in Diffusion Models

### v2

@@ 3.1 Experimental Results
On CIFAR-10, we observed that the SFD-Two Stage model (or Two Stage, for short), which involves first distilling the pre-trained diffusion model with 50,000 steps and then fine-tuning it using the SFD loss for the same number of steps, exhibited faster forgetting. In Figure 7, we report two performance metrics, FID and UA, during the unlearning stage, compared with the results from SFD. The results indicate that SFD consistently outperforms the two-stage approach in both metrics given sufficient training. Although the two-stage approach started with a lower FID than SFD, its performance fluctuated and declined over time. The UA initially increased rapidly, peaked, and then slightly decreased at the end. The gain in UA during the unlearning stage came at the cost of FID. In contrast, SFD effectively coordinated machine unlearning and distillation to forget specific classes while retaining the original generative capability for the remaining classes, thereby improving both FID and UA throughout finetuning and achieving better final results. Nonetheless, the two-stage approach remains practical, especially when forgetting requirements vary over time or when there is an urgent need, as it appears more flexible and efficient under such conditions.
@@
We provide both qualitative and quantitative results of celebrity forgetting tasks on two selected celebrities, $i.e.$, Brad Pitt and Angelina Jolie, where the concepts to forget are “bard pitt” and “angelina jolie”, respectively, and the corresponding concepts to override are “a middle aged man” and “a middle aged woman”, respectively. As is shown in Figure 1 and Table 2, we showcase the effectiveness of SFD for forgetting certain concepts in text-to-image diffusion models, such as removing the generative capability of celebrities.
@@
SD v1.4 (Rombach et al., 2022) 10.4% 60.6% 11.7% 73.8%
SLD Medium (Schramowski et al., 2023) 14.1% 0.47% 11.9% 3.29%
ESD-x (Gandikota et al., 2023) 34.7% 2.01% 32.6% 3.35%
SA (Heng and Soh, 2024) 5.8% 7.52% 4.4% 7.74%
@@
In addition to the celebrity forgetting experiments, we conducted experiments on a broader concept forgetting task, namely, forgetting “nudity” as a concept. We note that “nudity” is a broader concept than individuals ($e.g.,$ celebrities) and forgetting “nudity” in general is much more challenging. Therefore, we adopted a slightly different strategy for this task to enhance the forgetting performance. In particular, we first created a list of 12 common human subjects (see Table 5) that can be potentially misused for generating “nudity”-related contents and randomly paired them with one of NSFW keywords (see Table 6) as prompts to forget. We further leveraged the negative prompting technique to match these prompts with their corresponding prompts to override. Specifically, we take the original text prompt as the conditional text input while using the concatenated NSFW keywords instead of an empty string as the unconditional text input. We notice this approach also has a concept forgetting effect on the original score distillation method, which is denoted as “SiD-LSG-Neg.” We report key MU performance metrics in Table 3. Sample images by baselines and SFD are displayed in Figure 6.
@@
Model Inapprop. Prob. ($\downarrow$) Max. Exp. Inapprop. ($\downarrow$)
SD v1.4 (Rombach et al., 2022) 28.54% 86.6%
SiD-LSG (Zhou et al., 2024a) 26.86% 88.12%
SiD-LSG-Neg (Ours) 20.97% 81.64%
SLD Medium (Schramowski et al., 2023) 14.10% 71.73%
ESD-u (Gandikota et al., 2023) 16.94 % 69.68%
SFD-Two Stage (Ours) 11.03% 66.90%
Figure 6: Generated images using different text-to-image diffusion models. The prompts used for generation follow the general form, “A photo of a <nudity keyword> <human subject>.” Sensitive body parts were manually censored after image generation.
@@ 3.2 Ablation Studies
EDM (Karras et al., 2022) is a state-of-the-art diffusion model with enhanced capability for generating high-quality images. To evaluate our method’s generalizability across different model architectures, we additionally conduct experiment using the EDM architecture. We adapted the codebase used by SiD (Zhou et al., 2024b) and fine-tuned the pre-trained class-conditional CIFAR-10 EDM-VP model. Figure 5 shows that the FID results of our method can be further improve when based on a more powerful pre-trained model.
@@
Classifier-free guidance (CFG), first proposed by Ho and Salimans (2021), is a commonly-used strategy for conditional sampling. While typically adopted during inference to enhance class fidelity, it has also been shown to be useful for the training of score-based distillation (Yin et al., 2024; Zhou et al., 2024a). We compare our models trained with and without CFG in Table 4. In our experiments on STL-10, we found that including classifier-free guidance during training improved the performance in terms of both FID and UA. However, we did not observe such improvements on the CIFAR-10 dataset; on the contrary, we noticed a degradation in the evaluation metrics. We speculate that the influence of CFG may be tied to the inter-class differences: when training data contain classes sharing similar features, such as automobile and truck in CIFAR-10, training with CFG may not be as beneficial as it is when the training dataset consists of more distinct classes.
@@ 4 Conclusion
Our work demonstrates the benefits of the proposed score forgetting distillation (SFD), which achieves accelerated forgetting with score-based distillation, providing a unified and effective solution to diffusion-based generative modeling and machine unlearning. The generator trained by our method produces high-quality images of desired classes with a single step while the target class is effectively forgotten. Our experiments show that the proposed strategy attains noticeable gains in performance on both CIFAR-10 and STL-10. We further conduct the detailed study with the SFD in different settings, e.g., comparing SFD against baselines as well as different configurations of SFD in terms of UA, FID, and other metrics. Additionally, we provide qualitative results of concept forgetting for text-to-image diffusion models like SD. To summarize, the forgetting method is effective and general, with the potential to be incorporated into existing models, such as text-to-image diffusion models.
Appendix for Score Forgetting Distillation
@@ Appendix A Related work
The study of MU can be traced back to classical machine learning models in response to data protection regulations such as “the right to be forgotten” [Cao and Yang, 2015, Hoofnagle et al., 2019, Bourtoule et al., 2021, Nguyen et al., 2022]. Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification [Ginart et al., 2019, Golatkar et al., 2020, Neel et al., 2021, Sekhari et al., 2021], text-to-image generation [Gandikota et al., 2023, Zhang et al., 2023a, Kumari et al., 2023, Fan et al., 2024], federated learning [Halimi et al., 2022, Che et al., 2023], and graph neural networks [Chen et al., 2022, Chien et al., 2022, Wu et al., 2023]. In the literature, ‘exact’ unlearning, which involves retraining the model from scratch after removing specific training data points, is often considered the gold standard. However, this approach comes with significant computational demands and requires access to the entire training set [Thudi et al., 2022]. To address these challenges, many research efforts have shifted towards the development of scalable and effective approximate unlearning methods [Liu et al., 2024, Chen et al., 2023a]. In addition, probabilistic methods with certain provable removal guarantees have been explored, often leveraging the concept of differential privacy [Neel et al., 2021, Sekhari et al., 2021]. Focusing on MU in diffusion-based image generation, this paper introduces a general data-free approach for rapid forgetting and one-step sampling in diffusion models, eliminating the need to access any real data.
@@
In examining the challenges and strategies associated with diffusion models and MU, several key issues and methodologies have been identified. Diffusion models, particularly when trained on data from open collections, face risks of contamination or manipulation, which could lead to the generation of inappropriate or offensive content [Chen et al., 2023b, Schramowski et al., 2023]. Strategies to mitigate these include data censoring and safety guidance to steer models away from undesirable outputs [Nichol et al., 2021], and introducing subtle perturbations to protect artistic styles [Shan et al., 2023]. Despite these measures, challenges remain in fully preventing diffusion models from generating harmful content or being susceptible to targeted poison attacks [Rando et al., 2022]. Furthermore, the evaluation of MU presents unique difficulties, especially as conventional retraining benchmarks are often impractical. Empirical metrics for assessing MU include unlearning accuracy, the utility of the model post-unlearning, and the use of classifiers to gauge the integrity of generated outputs [Jang et al., 2022]. Unlike existing methods, our approach efficiently suppresses the generation of harmful content using a one-step diffusion generator that overrides ‘unsafe’ concepts with MU-regularized score-based distillation.
@@
Generative modeling is a pivotal area in statistics and machine learning. Prior to the development of diffusion models and their associated denoising score matching (SM) techniques, effectively matching distributions in high-dimensional spaces—particularly those with intractable probability density functions—posed a significant challenge. Traditionally, deep generative models aimed to minimize discrepancies between data and model probability distributions using various distribution-matching related loss functions. These included Kullback-Leibler (KL) divergence [Kingma and Welling, 2013, Yin and Zhou, 2018], Jensen-Shannon (JS) divergence [Goodfellow et al., 2014], and transport cost [Tanwisuth et al., 2021, Zheng and Zhou, 2021, Zhang et al., 2021, Tanwisuth et al., 2023]. While VAEs and GANs developed under this framework have significantly advanced the field of generative modeling, they have exhibited limited capabilities in faithfully regenerating the original data. More recent methods have utilized data-based Fisher divergence [Song and Ermon, 2019, Ho et al., 2020, Song et al., 2020] to compare noise-corrupted data with noise-corrupted model distributions. While directly minimizing Fisher divergence, $i.e.$, the explicit SM loss, is intractable, diffusion models have effectively transformed the problem into minimizing a data-based denoising SM loss [Vincent, 2011, Sohl-Dickstein et al., 2015]. This transformation has allowed diffusion models to demonstrate exceptional capabilities in generating high-dimensional data that closely resemble the original distribution. However, the iterative denoising-based sampling inherent in these models is not only slow but also complicates efforts to further optimize the data generation process for downstream tasks. This issue becomes particularly challenging for tasks such as MU, which require the model to selectively forget specific concepts we are targeting in this paper.
@@
Classic score-matching-based diffusion models [Sohl-Dickstein et al., 2015, Song and Ermon, 2019, Ho et al., 2020, Song et al., 2020] have become increasingly influential in developing generative models with high extensibility and sample quality [Dhariwal and Nichol, 2021, Karras et al., 2022, Ramesh et al., 2022]. However, standard Gaussian diffusion models, along with other non-Gaussian variants [Hoogeboom et al., 2021, Austin et al., 2021, Chen and Zhou, 2023, Zhou et al., 2023], suffer from relatively slow sampling compared to traditional one-step generative models, such as GANs and VAEs. Inspired by the success of applying diffusion processes to the training of generative models, Xiao et al. [2021] and Wang et al. [2022] were among the first to promote faster generation by leveraging both adversarial training techniques and diffusion-based data augmentation. However, these approaches inevitably reintroduce potential issues like training instability and mode collapse. Closely related to the original score matching, Salimans and Ho [2022] proposed progressively halving the steps needed in the reverse generation process. Similarly, Song et al. [2023] presented the consistency model as a method for distilling the reverse ODE sampling process. Along this direction, much effort has been made by others [Xu et al., 2023, Yin et al., 2024, Luo et al., 2023, Zhou et al., 2024b] to improve both sample quality and diversity.
@@
To address the slow sampling speed associated with traditional diffusion models, score distillation methods have been developed to harness pretrained score functions. These methods approximate data scores, facilitating model distribution matching under noisy conditions to align with the noisy data distribution governed by the pretrained denoising score matching function. These methods, as explored in several recent works [Poole et al., 2023, Wang et al., 2023, Luo et al., 2023, Nguyen and Tran, 2023, Yin et al., 2024], primarily utilize the KL divergence, whose gradients can be analytically computed using both the pretrained and estimated score functions. Importantly, these KL-based methods do not require access to real data, as the KL divergence is defined with respect to the model distribution. While these approaches have successfully approximated the data distribution in a data-free manner, they often suffer from performance degradation when compared to the original, pretrained teacher diffusion model. Consequently, additional loss terms that require access to the original training data or data synthesized with the pretrained diffusion models are often necessary to mitigate this performance degradation. However, employing these terms voids the data-free feature of the process.
In response to these challenges, Score identity Distillation (SiD) has emerged as an effective data-free solution for matching distributions by minimizing a model-based Fisher divergence. Although directly computing this divergence is intractable, its minimization is effectively converted into a model-based score distillation loss. This data-free method facilitates the distillation of the pretrained score function from the teacher diffusion model into a potentially superior one-step student generator. Inspired by the success of this data-free score distillation, we are motivated to integrate its loss into our algorithm, SFD, to enhance its effectiveness and efficiency in generative modeling with data-free unlearning.
@@
When applying MU to classification tasks, effectiveness-oriented metrics include unlearning accuracy, which evaluates how accurately the model performs on the forget set after unlearning [Golatkar et al., 2020]. Utility-oriented metrics include remaining accuracy, which measures the updated model’s performance on the retain set post-unlearning [Song and Mittal, 2021], and testing accuracy, which assesses the model’s generalization capability after unlearning. For generation tasks, accuracy-based metrics use a post-generation classifier to evaluate the generated content [Zhang et al., 2023b], while quality metrics assess the overall utility of the generated outputs [Gandikota et al., 2023]. A significant limitation of these metrics, particularly in measuring unlearning effectiveness, is their heavy dependence on the specific unlearning tasks [Fan et al., 2024]. To address this, we train an external classifier to evaluate unlearning accuracy (UA), ensuring that the generated images do not belong to the forgetting class or concept. Additionally, we use FID to evaluate the quality of image generations for non-forgetting classes or prompts.

### v3

@@ 3.1 Experimental Results
On CIFAR-10, we observed that the SFD-Two Stage model (or Two Stage, for short), which involves first distilling the pre-trained diffusion model with 50,000 steps and then fine-tuning it using the SFD loss for the same number of steps, exhibited faster forgetting. In Figure 7, we report two performance metrics, FID and UA, during the unlearning stage, compared with the results from SFD. The results indicate that SFD consistently outperforms the two-stage approach in both metrics given sufficient training. Although the two-stage approach started with a lower FID than SFD, its performance fluctuated and declined over time. The UA initially increased rapidly, peaked, and then slightly decreased at the end. The gain in UA during the unlearning stage came at the cost of FID. In contrast, SFD effectively coordinated machine unlearning and distillation to forget specific classes while retaining the original generative capability for the remaining classes, thereby improving both FID and UA throughout finetuning and achieving better final results. Nonetheless, the two-stage approach remains practical, especially when forgetting requirements vary over time or when there is an urgent need, as it appears more flexible and efficient under such conditions. Specifically, with SFD-Two Stage, finetuning achieves more than a 10×10$\times$10 × speedup, delivering competitive results (FID = 5.73, UA = 99.5%) in as few as $\sim$1.5k steps.
@@
We provide both qualitative and quantitative results of celebrity forgetting tasks on two selected celebrities, $i.e.$, Brad Pitt and Angelina Jolie, where the concepts to forget are “brad pitt” and “angelina jolie,” respectively, and the corresponding concepts to override are “a middle aged man” and “a middle aged woman,” respectively. As shown in Figure 1 and Table 2, we showcase the effectiveness of SFD in forgetting concepts such as specific celebrities in text-to-image diffusion models. For this experiment, we exclude the previous baseline, SalUn, as the original paper did not evaluate its performance on the celebrity forgetting task.
@@
SD v1.4 [61] 10.4% 60.6% 11.7% 73.8%
SLD Medium [65] 14.1% 0.47% 11.9% 3.29%
ESD-x [20] 34.7% 2.01% 32.6% 3.35%
SA [27] 5.8% 7.52% 4.4% 7.74%
@@
In addition to the celebrity forgetting experiments, we conducted concept forgetting experiments for a broader concept, namely, “nudity.” We note that nudity is a more abstract and generalized concept than specific individuals (e.g., celebrities), making it a significantly more challenging forgetting task. To enhance the forgetting performance, we adopted a slightly different strategy for this task. In particular, we first curated a list of 12 common human subjects (see Table 5) that could potentially be misused for generating “nudity”-related content. Each subject was then randomly paired with one of the NSFW keywords (see Table 6) to create prompts to forget. Furthermore, we leveraged the negative prompting technique to associate these prompts with their corresponding prompts to override. Specifically, we used the original text prompt as the conditional text input while using the concatenated NSFW keywords instead of an empty string as the unconditional text input. We observed that this approach also induces a concept forgetting effect in the original score distillation method, which we denote as “SiD-LSG-Neg.” Key MU performance metrics are reported in Table 3, while sample images generated by baseline methods and SFD are displayed in Figure 6.
@@
Model Inapprop. Prob. ($\downarrow$) Max. Exp. Inapprop. ($\downarrow$) CLIP ($\uparrow$)
SD v1.4 [61] 28.54% 86.6% 31.93
SiD-LSG [95] 26.86% 88.12% 31.23
SiD-LSG-Neg (Ours) 20.97% 81.64% 31.22
SLD Medium [65] 14.10% 71.73% 30.77
ESD-u [20] 16.94 % 69.68% 30.15
SFD-Two Stage (Ours) 11.03% 66.90% 30.25
Figure 6: Generated images using different text-to-image diffusion models. The prompts used to generate are in a general form of “A photo of a <nudity keyword> <human subject>.” Sensitive body parts are manually censored after generation.
@@ 3.2 Ablation Studies
EDM [39] is a state-of-the-art diffusion model with enhanced capability for generating high-quality images. To evaluate our method’s generalizability across different model architectures, we additionally conduct experiment using the EDM architecture. We adapted the codebase used by SiD [96] and fine-tuned the pre-trained class-conditional CIFAR-10 EDM-VP model. Figure 5 shows that the FID results of our method can be further improve when based on a more powerful pre-trained model.
@@
Classifier-free guidance (CFG), first proposed by Ho & Salimans [29], is a commonly-used strategy for conditional sampling. While typically adopted during inference to enhance class fidelity, it has also been shown to be useful for the training of score-based distillation [86; 95]. We compare our models trained with and without CFG in Table 4. In our experiments on STL-10, we found that including classifier-free guidance during training improved the performance in terms of both FID and UA. However, we did not observe such improvements on the CIFAR-10 dataset; on the contrary, we noticed a degradation in the evaluation metrics. We speculate that the influence of CFG may be tied to the inter-class differences: when training data contain classes sharing similar features, such as automobile and truck in CIFAR-10, training with CFG may not be as beneficial as it is when the training dataset consists of more distinct classes.
@@ 4 Conclusion
Our work demonstrates that the proposed method, SFD, achieves accelerated forgetting through score-based distillation, providing an effective solution to diffusion-based generative modeling and MU. Specifically, the SFD model produces high-quality images in a single step, while ensuring that the target class or concept is effectively forgotten. Experiments on CIFAR-10 and STL-10 validate the effectiveness of our approach, showing notable improvements in performance. We further perform a detailed analysis of SFD across various settings, including comparisons against baselines and different configurations of SFD in terms of UA, FID, and other metrics. Additionally, we provide both qualitative and quantitative results on concept forgetting in text-to-image diffusion models such as Stable Diffusion. To summarize, SFD introduces a novel, data-free solution for MU in diffusion models, which also significantly accelerates their sampling speed.
Acknowledgments
The authors acknowledge the support of NSF-IIS 2212418, NIH-R37 CA271186, and a gift fund from Apple.
Appendix
@@ Appendix A Related work
The study of MU can be traced back to classical machine learning models in response to data protection regulations such as “the right to be forgotten” [6; 33; 5; 51]. Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification [21; 22; 50; 66], text-to-image generation [20; 42; 87; 19], federated learning [25; 7], and graph neural networks [8; 12; 81]. In the literature, ‘exact’ unlearning, which involves retraining the model from scratch after removing specific training data points, is often considered the gold standard. However, this approach comes with significant computational demands and requires access to the entire training set [76]. To address these challenges, many research efforts have shifted towards the development of scalable and effective approximate unlearning methods [44; 9]. In addition, probabilistic methods with certain provable removal guarantees have been explored, often leveraging the concept of differential privacy [50; 66]. Focusing on MU in diffusion-based image generation, this paper introduces a general data-free approach for rapid forgetting and one-step sampling in diffusion models, eliminating the need to access any real data.
@@
In examining the challenges and strategies associated with diffusion models and MU, several key issues and methodologies have been identified. Diffusion models, particularly when trained on data from open collections, face risks of contamination or manipulation, which could lead to the generation of inappropriate or offensive content [11; 65]. Strategies to mitigate these include data censoring and safety guidance to steer models away from undesirable outputs [53], and introducing subtle perturbations to protect artistic styles [67]. Despite these measures, challenges remain in fully preventing diffusion models from generating harmful content or being susceptible to targeted poison attacks [59]. Furthermore, the evaluation of MU presents unique difficulties, especially as conventional retraining benchmarks are often impractical. Empirical metrics for assessing MU include unlearning accuracy, the utility of the model post-unlearning, and the use of classifiers to gauge the integrity of generated outputs [37]. Unlike existing methods, our approach efficiently suppresses the generation of harmful content using a one-step diffusion generator that overrides ‘unsafe’ concepts with MU-regularized score-based distillation.
Concept Erasure for Diffusion Models
Diffusion models have gained significant attention and also triggered many controversies due to their incredible capability of generating high-quality, diverse visual content. For example, with ill-intended text prompts, text-to-image diffusion models can easily generate inappropriate images containing sensitive content. Consequently, concept erasure (CE) has become a high priority for mitigating such problems. Current approaches mainly fall into two categories: sampling-based training-free approaches and finetuning-based MU approaches. One classic sampling-based approach is to set concepts to erase as negative prompts during sampling, which is a direct application of classifier-free guidance (CFG) [29]. Further enhancing the idea of safe guidance, Schramowski et al. [65] propose Safe Latent Diffusion (SLD) as a configurable method to balance suppressing “unsafe” concepts with minimizing its impact on generated images. In parallel, finetuning-based MU methods have also been applied to solve concept erasure problems [20; 27; 87; 19]. Closely related to CFG, ESD [20] finetunes the Stable Diffusion components to fit a target conditional score function that contains the opposite direction of the score associated with concepts to remove. Heng & Soh [27] perceive the MU problem from a Bayesian continual learning perspective and introduce replaying data to retain the model’s generative capability for data to remember. Zhang et al. [87] present a cross-attention-based loss to tackle the problem by minimizing attention weights related to the concepts to forget. To improve finetuning efficiency, Fan et al. [19] propose selecting parameters for finetuning based on the saliency map of the concept to remove. However, existing methods are all based on standard multi-step diffusion models, making them not directly compatible with more efficient one-step diffusion models distilled using score distillation methods. Therefore, we foresee an opportunity for a novel, swift, and data-free MU approach that leverages score distillation to solve the data-free MU problem while simultaneously enhancing the distilled model’s resilience to "unsafe" concepts, achieving both goals at once.
@@
Generative modeling is a pivotal area in statistics and machine learning. Prior to the development of diffusion models and their associated denoising score matching (SM) techniques, effectively matching distributions in high-dimensional spaces—particularly those with intractable probability density functions—posed a significant challenge. Traditionally, deep generative models aimed to minimize discrepancies between data and model probability distributions using various distribution-matching related loss functions. These included Kullback-Leibler (KL) divergence [40; 85], Jensen-Shannon (JS) divergence [23], and transport cost [74; 92; 88; 75]. While VAEs and GANs developed under this framework have significantly advanced the field of generative modeling, they have exhibited limited capabilities in faithfully regenerating the original data. More recent methods have utilized data-based Fisher divergence [71; 31; 72] to compare noise-corrupted data with noise-corrupted model distributions. While directly minimizing Fisher divergence, $i.e.$, the explicit SM loss, is intractable, diffusion models have effectively transformed the problem into minimizing a data-based denoising SM loss [77; 69]. This transformation has allowed diffusion models to demonstrate exceptional capabilities in generating high-dimensional data that closely resemble the original distribution. However, the iterative denoising-based sampling inherent in these models is not only slow but also complicates efforts to further optimize the data generation process for downstream tasks. This issue becomes particularly challenging for tasks such as MU, which require the model to selectively forget specific concepts we are targeting in this paper.
@@
Classic score-matching-based diffusion models [69; 71; 31; 72] have become increasingly influential in developing generative models with high extensibility and sample quality [16; 39; 58]. However, standard Gaussian diffusion models, along with other non-Gaussian variants [34; 3; 10; 94], suffer from relatively slow sampling compared to traditional one-step generative models, such as GANs and VAEs. Inspired by the success of applying diffusion processes to the training of generative models, Xiao et al. [82] and Wang et al. [78] were among the first to promote faster generation by leveraging both adversarial training techniques and diffusion-based data augmentation. However, these approaches inevitably reintroduce potential issues like training instability and mode collapse. Closely related to the original score matching, Salimans & Ho [63] proposed progressively halving the steps needed in the reverse generation process. Similarly, Song et al. [73] presented the consistency model as a method for distilling the reverse ODE sampling process. Along this direction, much effort has been made by others [83; 86; 47; 96] to improve both sample quality and diversity.
@@
To address the slow sampling speed associated with traditional diffusion models, score distillation methods have been developed to harness pretrained score functions. These methods approximate data scores, facilitating model distribution matching under noisy conditions to align with the noisy data distribution governed by the pretrained denoising score matching function. These methods, as explored in several recent works [57; 79; 47; 52; 86], primarily utilize the KL divergence, whose gradients can be analytically computed using both the pretrained and estimated score functions. Importantly, these KL-based methods do not require access to real data, as the KL divergence is defined with respect to the model distribution. While these approaches have successfully approximated the data distribution in a data-free manner, they often suffer from performance degradation when compared to the original, pretrained teacher diffusion model. Consequently, additional loss terms that require access to the original training data or data synthesized with the pretrained diffusion models are often necessary to mitigate this performance degradation. However, employing these terms voids the data-free feature of the process. In response to these challenges, Score identity Distillation (SiD) has emerged as an effective data-free solution for matching distributions by minimizing a model-based Fisher divergence. Although directly computing this divergence is intractable, its minimization is effectively converted into a model-based score distillation loss. This data-free method facilitates the distillation of the pretrained score function from the teacher diffusion model into a potentially superior one-step student generator. Inspired by the success of this data-free score distillation, we are motivated to integrate its loss into our algorithm, SFD, to enhance its effectiveness and efficiency in generative modeling with data-free unlearning.
@@
When applying MU to classification tasks, effectiveness-oriented metrics include unlearning accuracy, which evaluates how accurately the model performs on the forget set after unlearning [22]. Utility-oriented metrics include remaining accuracy, which measures the updated model’s performance on the retain set post-unlearning [70], and testing accuracy, which assesses the model’s generalization capability after unlearning. For generation tasks, accuracy-based metrics use a post-generation classifier to evaluate the generated content [89], while quality metrics assess the overall utility of the generated outputs [20]. A significant limitation of these metrics, particularly in measuring unlearning effectiveness, is their heavy dependence on the specific unlearning tasks [19]. To address this, we train an external classifier to evaluate unlearning accuracy (UA), ensuring that the generated images do not belong to the forgetting class or concept. Additionally, we use FID to evaluate the quality of image generations for non-forgetting classes or prompts.